OpenStudy (anonymous):

Why McLaurin of f(x)= e^(cos(x^3)) is e - (ex^6)/6? Why there's e?

OpenStudy (anonymous):

because is the equation of E=1

OpenStudy (anonymous):

the equation of e =1? Sorry but I don't understand what you mean.

OpenStudy (anonymous):

NO ,, because the E is equal to 1 the answer is (cosx1(e)

myininaya (myininaya):

the series goes like f(x)=f(0)+f'(0)*x+f''(0)*x^2/2!+f'''(0)*x^3/3!+..... so f(0)=e^(cos0^3)=e^(cos0)=e^(1)=e

myininaya (myininaya):

cos0=1

myininaya (myininaya):

hector do you understand why the constant e is there now?

OpenStudy (anonymous):

oh that's clear. Thanks you.

myininaya (myininaya):

thanks :)

OpenStudy (anonymous):

what